Differentiate f(x) = (8x3 + 4x )(2x2 +5) using product rule.
1
Expert's answer
2020-09-10T19:05:09-0400
"Solution"

"f(x)=h(x) \cdot g(x)\\
f'(x)=h'(x) \cdot g(x)+h(x) \cdot g'(x)\\
f(x)=(8x^3+4x)(2x^2+5)\\
f'(x)=(24x^2+4)(2x^2+5)+(8x^3+4x)(4x)\\
Expanding\\
=48x^4+120x^2+8x^2+20+32x^4+16x^2\\

\implies f'(x)=80x^4+144x^2+20-------->Answer"